OpenStudy (anonymous):

A = (f+g/2) h for g SHOW WORK PLEASE

OpenStudy (anonymous):

A = (f+g/2) h for g \[a \div h=(f+g/2)\] \[a \div h\]= f+g/2 (\[2\times a \div h=f+g\] \[2\times a \div h-g=f\]